# Mildly Super-Eddington Accretion Onto Slowly-Spinning Black Holes Explains the X-Ray Weakness of the Little Red Dots

### v1

@@ 5 Summary and Conclusions
This study was motivated by the widespread JWST detection at $z>4$ of sources characterized by solid evidence for the presence of low-luminosity AGN, with SMBH masses of $M_{\bullet}=10^{6-8}\,{\rm M_{\odot}}$. These SMBHs, found in compact and red hosts (the LRDs), are largely undetected in X-rays, presenting the “X-ray weakness” problem.
Using a suite of GRRMHD simulations, we investigated the super-Eddington accretion process onto a SMBH with mass $M_{\bullet}=10^{7}\,{\rm M_{\odot}}$ at $z\sim 6$; these values represent the medians for the LRDs population in the largest catalogs thus far available (Harikane et al., 2023; Maiolino et al., 2023; Kocevski et al., 2024; Akins et al., 2024). Our key findings are summarized here.
@@
  • About $55\%$ of our super-Eddington SEDs have optical-UV to X-ray ratios ($\,{\alpha_{\rm ox}}$) outside the typical range for standard Type 1 AGN. Contrary to the standard, our SEDs show increasing X-ray weakness for increasing optical-UV luminosity.
@@
  • The effect of the viewing angle, or beaming, is crucial. The same SMBH, with the same Eddington ratio and spin, produces SEDs whose peaks differ by $1.5$ dex in flux when the viewing angle changes from $80^{\circ}$ to $10^{\circ}$. In general, it is most likely to observe such systems along a line of sight that leads to an X-ray weak SED. Observers aligned with the jet axis would predominantly see X-rays. Conversely, observers at larger angles see cooler regions, resulting in spectra dominated by optical-UV emission.
  • The inner structure of the accretion flow reveals the physical difference between X-ray weak and strong SEDs. Enhanced X-ray emission is linked to higher temperatures ($T>10^{8}$ K) close to the SMBH and a highly relativistic jet, with higher spin values leading to more efficient jet production and harder X-ray spectra.
@@
Acknowledgments: F.P. acknowledges fruitful discussions with Andrea Ferrara, Erin Kara, Roberto Maiolino, Nico Cappelluti, and Minghao Yue. F.P. also acknowledges support from a Clay Fellowship administered by the Smithsonian Astrophysical Observatory. This work was also supported by the Black Hole Initiative at Harvard University, funded by grants from the John Templeton Foundation and the Gordon and Betty Moore Foundation. This work used the Anvil supercomputer at Purdue University through allocation AST-080028 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296.

### v2

@@ 5 Summary and Conclusions
This study was motivated by the widespread JWST detection at $z>4$ of sources characterized by solid evidence for the presence of low-luminosity AGN, with SMBH masses of $M_{\bullet}=10^{6-8}\,{\rm M_{\odot}}$. These SMBHs, found in compact and red hosts (the LRDs), are largely undetected in X-rays, leading to the “X-ray weakness” problem.
Using a suite of GRRMHD simulations, we investigated the super-Eddington accretion process onto a SMBH with mass $M_{\bullet}=10^{7}\,{\rm M_{\odot}}$ at $z\sim 6$; these values represent the medians for the LRDs population in the largest catalogs thus far available (Harikane et al., 2023; Maiolino et al., 2023; Kocevski et al., 2024; Akins et al., 2024; Kokorev et al., 2024). Our key findings are summarized here.
@@
  • About $55\%$ of our super-Eddington SEDs have optical-UV to X-ray ratios ($\,{\alpha_{\rm ox}}$) outside the typical range for standard Type 1 AGN. Contrary to the standard, our SEDs show increasing X-ray emission for increasing optical-UV luminosity.
@@
  • The effect of the viewing angle, or beaming, is crucial. The same SMBH, with the same Eddington ratio and spin, produces SEDs whose peaks differ by $1.5$ dex in flux when the viewing angle changes from $80^{\circ}$ to $10^{\circ}$. In general, observers aligned with the jet axis are most likely to see X-rays. Observers at larger and more probable viewing angles see cooler regions, resulting in spectra dominated by optical-UV emission.
  • The inner structure of the accretion flow reveals the physical difference between X-ray weak and strong SEDs. Enhanced X-ray emission is linked to higher temperatures ($T>10^{8}$ K) close to the SMBH and a highly relativistic jet, with higher spin values leading to more powerful jets and harder X-ray spectra.
@@
  • The phenomenon of the LRDs is mainly observed in the redshift range $84<z<$ 8, where large availability of accretable gas would render widespread super-Eddington accretion feasible. An older Universe, with far less free gas available, would make this phenomenon disappear, as observed.
@@
Acknowledgments: F.P. acknowledges fruitful discussions with Nico Cappelluti, Andrea Ferrara, Erin Kara, Vasily Kokorev, Roberto Maiolino, Joseph Silk, and Minghao Yue. F.P. also acknowledges support from a Clay Fellowship administered by the Smithsonian Astrophysical Observatory. This work was also supported by the Black Hole Initiative at Harvard University, funded by grants from the John Templeton Foundation and the Gordon and Betty Moore Foundation. This work used the Anvil supercomputer at Purdue University through allocation AST-080028 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296.